OpenStudy (anonymous):

multiply (z+3)(z^2+2z-3)

OpenStudy (anonymous):

\[\tt{Foil\space this\space out}\]

OpenStudy (anonymous):

how can you foil?

OpenStudy (anonymous):

\[z*z^2+2z*z+z*-3+3*z^2+3*2z+3*-3\]

OpenStudy (anonymous):

would it be z^3+6z^2 -18z-9

OpenStudy (anonymous):

I got -9 + 3 z + 5 z^2 + z^3

OpenStudy (anonymous):

Multiply everything in the second bracket by z, then multiply everything in the second bracket by 3. Group like items together and add.

OpenStudy (anonymous):

gotcha

OpenStudy (anonymous):

F - First O - Outer I - Inner L - Last Ex: =(x+y)(a+b) =ax + bx + ay + by